# FlowSE: Efficient and High-Quality Speech Enhancement via Flow Matching

### v1

@@ Abstract
Generative models have excelled in audio tasks using approaches such as language models, diffusion, and flow matching. However, existing generative approaches for speech enhancement (SE) face notable challenges: language model-based methods suffer from quantization loss, leading to compromised speaker similarity and intelligibility, while diffusion models require complex training and high inference latency. To address these challenges, we propose FlowSE, a flow-matching-based model for SE. Flow matching learns a continuous transformation between noisy and clean speech distributions in a single pass, significantly reducing inference latency while maintaining high-quality reconstruction. Specifically, FlowSE trains on noisy mel spectrograms and optional character sequences, optimizing a condition flow matching loss with ground-truth mel spectrograms as supervision. It implicitly learns speech’s temporal-spectral structure and text-speech alignment. During inference, FlowSE can operate with or without textual information, achieving impressive results in both scenarios, with further improvements when transcripts are available. Extensive experiments demonstrate that FlowSE significantly outperforms state-of-the-art generative methods, establishing a new paradigm for generative-based SE and demonstrating the potential of flow matching to advance the field. Our code, pre-trained checkpoints, and audio samples are available [^1] https://github.com/Honee-W/FlowSE/.
@@ 1 Introduction
Our contributions are summarized as follows: (1) We introduce FlowSE, a rectified flow matching framework with a DiT backbone for efficient, high-fidelity SE that preserves speaker identity. (2) Extensive experiments demonstrate that FlowSE significantly improves speech quality, intelligibility, and speaker similarity while reducing inference latency compared to existing generative SE methods.
@@ 4.1 Datasets & Evaluation Metrics
Training sets To comprehensively evaluate the effectiveness of FlowSE, we construct a large-scale training dataset by combining multiple publicly available speech and noise datasets. We utilize WeNetSpeech [25], GigaSpeech [26], VoiceBank(VCTK) [27], and the DNS Challenge - Interspeech 2021 datasets [28] to provide speech recordings. We incorporate environmental and artificial noise from the WHAM! [29], DEMAND [30] and the DNS Challenge - Interspeech 2021 datasets [28]. We use room impulse responses from the OpenSLR26 and OpenSLR28 datasets [31]. Non-English and non-Chinese utterances are filtered out. Transcriptions are obtained using OpenAI’s Whisper model [32] [^2] https://huggingface.co/openai/whisper-large-v3-turbo.
@@
Table 1: Comparison of different systems on the DNS Challenge test set. Higher values indicate better performance.
@@ 4.2 Training Setup
We use 100-dimensional log mel-filterbank features extracted from 24 kHz audio with a hop length of 256 as input representation. We adopt a latent Diffusion Transformer (DiT) [22] as our velocity field estimator. The model consists of N = 22 layers, 16 attention heads, a hidden dimension of 1024, and a feed-forward network (FFN) dimension of 2048. We utilize ConvNeXt V2 [35] as the text encoder, with an embedding dimension of 512 and an FFN dimension of 1024. It is trained to process character sequences extracted from transcriptions. A pre-trained Vocos vocoder [23] is used to synthesize waveforms from the enhanced mel spectrograms.
We use the AdamW optimizer with a peak learning rate of $7.5\times 10^{-5}$, which is linearly warmed up over 20K steps and then linearly decayed. A dropout rate of 0.1 is applied to attention layers and FFN layers. The gradient norm is clipped to a maximum value of 1 to stabilize training. The model is trained on 8 NVIDIA 4090D GPUs for 1.2 million updates.

### v2

@@ Abstract
Generative models have excelled in audio tasks using approaches such as language models, diffusion, and flow matching. However, existing generative approaches for speech enhancement (SE) face notable challenges: language model-based methods suffer from quantization loss, leading to compromised speaker similarity and intelligibility, while diffusion models require complex training and high inference latency. To address these challenges, we propose FlowSE, a flow-matching-based model for SE. Flow matching learns a continuous transformation between noisy and clean speech distributions in a single pass, significantly reducing inference latency while maintaining high-quality reconstruction. Specifically, FlowSE trains on noisy mel spectrograms and optional character sequences, optimizing a condition flow matching loss with ground-truth mel spectrograms as supervision. It implicitly learns speech’s temporal-spectral structure and text-speech alignment. During inference, FlowSE can operate with or without textual information, achieving impressive results in both scenarios, with further improvements when transcripts are available. Extensive experiments demonstrate that FlowSE significantly outperforms state-of-the-art generative methods, establishing a new paradigm for generative-based SE and demonstrating the potential of flow matching to advance the field. Our code, pre-trained checkpoints, and audio samples are available at https://github.com/Honee-W/FlowSE/.
@@
22footnotetext: Equal contribution.   11footnotemark: 1Corresponding author.
@@ 1 Introduction
In summary, we introduce FlowSE, a rectified flow matching framework with a DiT backbone for efficient, high-fidelity SE that preserves speaker identity. Extensive experiments demonstrate that FlowSE significantly improves speech quality, intelligibility, and speaker similarity while reducing inference latency compared to existing generative SE methods.
@@ 4.1 Datasets & Evaluation Metrics
Training sets To comprehensively evaluate the effectiveness of FlowSE, we construct a large-scale training dataset by combining multiple publicly available speech and noise datasets. We utilize WeNetSpeech [25], GigaSpeech [26], VoiceBank(VCTK) [27], and the DNS Challenge - Interspeech 2021 datasets [28] to provide speech recordings. We incorporate environmental and artificial noise from the WHAM! [29], DEMAND [30] and the DNS Challenge - Interspeech 2021 datasets [28]. We use room impulse responses from the OpenSLR26 and OpenSLR28 datasets [31]. Non-English and non-Chinese utterances are filtered out. Transcriptions are obtained using OpenAI’s Whisper model [32]***https://huggingface.co/openai/whisper-large-v3-turbo.
@@
Table 1: Comparison of different systems on the DNS Challenge test set.
@@ 4.2 Training Setup
We use 100-dimensional log mel-filterbank features extracted from 24 kHz audio with a hop length of 256 as input representation. We adopt a latent Diffusion Transformer (DiT) [22] as our velocity field estimator. The model consists of N = 22 layers, 16 attention heads, a hidden dimension of 1024, and a feed-forward network (FFN) dimension of 2048. We utilize ConvNeXt V2 [35] as the text encoder, with an embedding dimension of 512 and an FFN dimension of 1024. It is trained to process character sequences extracted from transcriptions. A pre-trained Vocos vocoder [23] is used to synthesize waveforms. We use the AdamW optimizer with a peak learning rate of $7.5\times 10^{-5}$, which is linearly warmed up over 20K steps and then linearly decayed. A dropout rate of 0.1 is applied to attention layers and FFN layers. The gradient norm is clipped to a maximum value of 1 to stabilize training. The model is trained on 8 NVIDIA 4090D GPUs for 1.2 million updates.